# Analysis of the Picard-Newton iteration for the incompressible Boussinesq equations

### v1

@@ 4.1 Differentially Heated Cavity
4.2 Boussinesq on a Complex Domain
Solutions for $u$ and $T$ shown on the domain in figure 4 for $Ra=1000$. We use a Delaunay mesh with max element diameter $h=7/64$, and again note that all results are comparable for other mesh sizes that we tested.
Figure 4: Shown above is the computed Boussinesq solution for $u$ (left) and $T$ (right) for $Ra=1000$ in the second numerical test.
Convergence results for Picard-Newton are shown in figure 5, for $Ra=5000,10000,15000,20000,25000,\text{ and }50000$. As in the previous test problem, we observe quadratic convergence for each $Ra$ which Picard-Newton converges, for (up to $Ra=10000$). For $Ra>10000$, we do not observe convergence within 200 iterations but note the computed solutions remains stable.
@@ 4.2 Boussinesq on a Complex Domain
Picard-Newton $20000$
Picard $1000$
Newton $5000$
Table 3: Shown above is the maximum $Ra$ (to the nearest 100) at which each method converges.
4.3 AAPicard-Newton
Anderson Acceleration (AA) is an extrapolation technique which utilizes previously computed iterates to construct the next solution. It is known to improve convergence for linearly convergent fixed point iterations like Picard [19, 7, 17, 22, 10, 7, 20]. Therefore, it is a natural addition to the Picard step of Picard-Newton and we call this iteration AAPicard-Newton.
4.3.1 Anderson-Acceleration
Algorithm 4.1.
AA for depth $m\geq 0$
Step 0: Choose $x_{0}\in Y$
Step 1: Find $w_{1}\in Y$ such that $w_{1}=g(x_{0})-x_{0}$. Set $x_{1}=x_{0}+w_{1}$.
Step k: For $k=2,3,...$. Set $m_{k}=\min\{k-1,m\}$.
  [a.] Find $w_{k}=g(x_{k-1})-x_{k-1}$.
  [b.] Solve the minimization problem for $\{\alpha_{j}^{k}\}_{k-m_{k}}^{k-1}$
  [c.] Set
where $w_{j}=g(x_{j-1})-x_{j-1}$ may be referred to as the update step or also as the nonlinear residual.
Setting $m=0$ returns the original fixed point iteration.
4.3.2 AAPicard-Newton for Differentially heated Cavity
We first apply AAPicard-Newton to the differentially heated cavity problem in section 4.1. Recall that for this problem Picard-Newton converges for $Ra$ as high as $250000$. Convergence for AAPicard-Newton, as shown in figure 7 for varying $Ra$ and $m=1$ (left) and $m=3$ (right). AAPicard-Newton shows significant improvement in convergence for higher $Ra$ both in iteration count and solvability. Convergence is even further improved with increased depth: for $m=1$ we observe convergence up to $Ra=500000$ and with $m=3$ convergence is obtained up to $Ra=750000$ (and convergence is accelerated by $m=3$ over $m=1$ for $Ra$ where $m=1$ converges).
Method Max $Ra$ for convergence
AAPicard-Newton: m=3 $750000$
AAPicard-Newton: m=1 $500000$
@@ 4.3.2 AAPicard-Newton for Differentially heated Cavity
Figure 7: Convergence plots for AAPicard-Newton with $m=1$ (left) and $m=3$ (right) at varying $Ra$.
4.3.3 AAPicard-Newton on a Complex Domain
We next apply AAPicard-Newton to the heated cavity problem on a complex domain from section 4.2. Recall that for this problem Picard-Newton converges in 200 iterations for $Ra$ up to $10000$. AAPicard-Newton achieves convergence with depth $m=1$ for $Ra$ up to $125000$ as shown in figure 8. This is further improved with depth $m=3$ where AAPicard-Newton converges for $Ra$ up to $75000$. As before, for higher $Ra$ we see stability for the method and it may converge if allowed to continue past 200 iterations.
@@ 4.3.3 AAPicard-Newton on a Complex Domain
AAPicard-Newton: m=3 $75000$
AAPicard-Newton: m=1 $125000$
@@
Figure 8: Shown above are convergence plots for AAPicard-Newton convergence with $m=1$ (left) and $m=3$ (right) at varying $Ra$.
@@ 5 Conclusion
The Picard-Newton iteration for the Boussinesq equations is an improvement upon the Newton iteration that provides less restrictive (sufficient) conditions for quadratic convergence. The sufficient conditions of Picard-Newton incorporate the accuracy bounds of the Picard step, which leads to a scaling of the sufficient conditions of Newton by a constant less than $1$. Furthermore, the sufficient conditions of Picard-Newton depend solely upon the fluid velocity $u^{k}$ due to the Picard step causing accuracy of $T^{k}$ to depend on $u^{k}$. This is in stark contrast to the usual Newton whose sufficient conditions depends upon both $u^{k}$ and $T^{k}$. Furthermore, Picard-Newton is stable for any $Ra$ and any initial guess.
These behaviors are reflected in the numerical tests where for a simple domain we see convergence for high $Ra$ and errors that remain stable. In comparison, Newton diverges for relatively low $Ra$ and Picard converges slower for the same $Ra$. This is further shown in the Boussinesq equations on a complex domain which is comparatively harder. Picard-Newton converges for high $Ra$ and is stable, while Picard and Newton either converges slower or fail to converge for these same $Ra$.
Lastly, with the popularity of AA, linearly convergent methods like Picard are becoming faster and more applicable to problems where it is usually slow (or unable) to converge. This beneficial effect carries over to Picard-Newton: AAPicard-Newton improved Picard-Newton by lowering the iteration count for which the method converges for $Ra$, and also allows convergence for higher $Ra$.
@@ 6 Acknowledgements
All authors were partially supported by NSF grant DMS 2011490.

### v2

@@ 4.1.1 Picard-Newton
4.1.2 AA-Picard-Newton for Differentially Heated Cavity
We now apply AA-Picard-Newton to the differentially heated cavity problem. Recall that for this problem Picard-Newton converges for $Ra$ as high as $250000$. Convergence for AA-Picard-Newton are shown in figure 4 for varying $Ra$ and $m=1$ (left) and $m=3$ (right). AA-Picard-Newton shows significant improvement in convergence for higher $Ra$ both in iteration count and solvability. Convergence is even further improved with increased depth: for $m=1$ we observe convergence up to $Ra=500000$ and with $m=3$ convergence is obtained up to $Ra=750000$ (and convergence is accelerated by $m=3$ over $m=1$ for $Ra$ where $m=1$ converges).
@@ 4.1.2 AA-Picard-Newton for Differentially Heated Cavity
AA-Picard-Newton: m=3 $750000$
AA-Picard-Newton: m=1 $500000$
@@
Figure 4: Convergence plots for AA-Picard-Newton with $m=1$ (left) and $m=3$ (right) at varying $Ra$.
4.2 Heated Cavity on a Complex Domain
Solutions for $u$ and $T$ shown on the domain in figure 5 for $Ra=1000$. We use a Delaunay mesh with max element diameter $h=7/64$, and again note that all results are comparable for other mesh sizes that we tested. We again begin the study of the convergence of (AA-)Picard-Newton for this heated cavity problem by considering Picard-Newton and then AA-Picard-Newton.
Figure 5: Shown above is the computed solution for $u$ (left) and $T$ (right) for $Ra=1000$ in the second numerical test.
4.2.1 Picard-Newton
The convergence results are shown in figure 6, for $Ra$ $=$ $5000$, $10000$, $15000$, $20000$, $25000$, and $50000$. As in the previous test problem, we observe quadratic convergence for each $Ra$ which Picard-Newton converges, for (up to $Ra=10000$). For $Ra>10000$, we do not observe convergence within 200 iterations but note the computed solutions remain stable.
4.2.2 AA-Picard-Newton on a Complex Domain
We next apply AA-Picard-Newton to the heated cavity problem on a complex domain from section 4.2. Recall that for this problem (AA-)Picard-Newton converges in 200 iterations for $Ra$ up to $10000$. AA-Picard-Newton achieves convergence with depth $m=1$ for $Ra$ up to $125000$ as shown in figure 8. This is further improved with depth $m=3$ where AA-Picard-Newton converges for $Ra$ up to $75000$. As before, for higher $Ra$ we see stability for the method and it may converge if allowed to continue past 200 iterations.
@@ 4.2.2 AA-Picard-Newton on a Complex Domain
AA-Picard-Newton: m=3 $75000$
AA-Picard-Newton: m=1 $125000$
@@
Figure 8: Shown above are convergence plots for AA-Picard-Newton convergence with $m=1$ (left) and $m=3$ (right) at varying $Ra$.
@@ 5 Conclusion
(AA-)Picard-Newton for the non-isothermal flow system is an improvement upon the Newton iteration that provides less restrictive (sufficient) conditions for quadratic convergence, and dramatically better convergence results on the benchmark problems. The sufficient conditions of Picard-Newton incorporate the accuracy bounds of the Picard step, which leads to a scaling of the sufficient conditions of Newton by a constant less than $1$. Furthermore, the sufficient conditions of Picard-Newton depend solely upon the fluid velocity $u^{k}$ due to the preconditioning by Picard causing accuracy of $T^{k}$ to depend on $u^{k}$. This is in stark contrast to the usual Newton whose sufficient conditions depends upon both $u^{k}$ and $T^{k}$. Furthermore, Picard-Newton is stable for any initial guess, provided the data satisfy a smallness condition that also implies uniqueness of the PDE system. These analytical results are extendable to AA-Picard-Newton with extra technical difficulties.
This improved theory is reflected in the numerical tests where for a simple domain we see convergence for high $Ra$ and errors that remain stable. In comparison, Newton diverges for relatively low $Ra$ and Picard converges slower for the same $Ra$. This is further shown in the non-isothermal flow system on a complex domain which is comparatively harder. (AA-)Picard-Newton converges for high $Ra$ and is proved stable, while Picard and Newton either converges slower or fail to converge for these same $Ra$.
@@ 6 Acknowledgements
The author was partially supported by NSF grant DMS 2011490.